# LoFi: Scalable Local Image Reconstruction with Implicit Neural Representation

### v1

@@ Abstract
Neural fields or implicit neural representations (INRs) have attracted significant attention in machine learning and signal processing due to their efficient continuous representation of images and 3D volumes. In this work, we build on INRs and introduce a coordinate-based local processing framework for solving imaging inverse problems, termed LoFi (Local Field). Unlike conventional methods for image reconstruction, LoFi processes local information at each coordinate separately by multi-layer perceptrons (MLPs), recovering the object at that specific coordinate. Similar to INRs, LoFi can recover images at any continuous coordinate, enabling image reconstruction at multiple resolutions. With comparable or better performance than standard CNNs for image reconstruction, LoFi achieves excellent generalization to out-of-distribution data and memory usage almost independent of image resolution. Remarkably, training on $1024\times 1024$ images requires just 3GB of memory—over 20 times less than the memory typically needed by standard CNNs. Additionally, LoFi’s local design allows it to train on extremely small datasets with less than 10 samples, without overfitting or the need for regularization or early stopping. Finally, we use LoFi as a denoising prior in a plug-and-play framework for solving general inverse problems to benefit from its continuous image representation and strong generalization. Although trained on low-resolution images, LoFi can be used as a low-dimensional prior to solve inverse problems at any resolution. We validate our framework across a variety of imaging modalities, from low-dose computed tomography to radio interferometric imaging.
@@ I Introduction
Imaging inverse problems are ubiquitous in domains like medicine (Wang et al., 2008), material science (Holler et al., 2017), remote sensing (Blahut, 2004) and cosmology (Kaiser and Squires, 1993). Deep learning has been, for a long time, the method of choice for solving inverse problems. In particular, convolutional neural networks (CNNs) like U-Net (Ronneberger et al., 2015) have shown remarkable performance across diverse tasks including computed tomography (CT) (Jin et al., 2017), magnetic resonance imaging (MRI) (Hyun et al., 2018), photoacoustic imaging (Davoudi et al., 2019) and inverse scattering (Wei and Chen, 2018). The success of the U-Net is primarily attributed to its multiscale architecture with a large receptive field that extracts features from the input image at different scales (Liu et al., 2022).
@@
Despite their strong performance on low-dimensional 2D imaging problems, deep learning architectures can become computationally expensive for high-dimensional images (Zhang et al., 2017a; Liang et al., 2021; Wang et al., 2021). This inefficiency arises because current deep neural networks reconstruct the entire image simultaneously, necessitating substantial memory for back-propagation during training. Additionally, interpreting such complex architectures (Zhang et al., 2021; Liang et al., 2021; Fabian et al., 2022) and analyzing their reconstructions can be challenging. Simplified neural architectures can enhance our understanding of reconstruction mechanisms, enabling the design of robust models with strong generalization (Aggarwal et al., 2018).
Recently, implicit neural representations (INRs) (Sitzmann et al., 2020; Atzmon and Lipman, 2020; Chabra et al., 2020) have emerged as a promising tool for representing continuous signals, images, and 3D volumes. Unlike most existing deep learning models that treat signals as discrete arrays, INRs map coordinates to signal values using a deep neural network, typically a multi-layer perceptron (MLP), resulting in a continuous signal representation. INRs have been widely applied to various tasks, including solving partial differential equations (PDEs) (Sitzmann et al., 2020; Vlašić et al., 2022; Khorashadizadeh et al., 2023a), 3D shape modeling (Chen and Zhang, 2019; Peng et al., 2020; Jiang et al., 2020), and image super-resolution (Chen et al., 2021b; Khorashadizadeh et al., 2023a). One of the most well-known applications of INRs is 3D scene representation using neural radiance fields (NeRF) (Mildenhall et al., 2021).
@@
INRs have several advantages over standard deep learning models. Rather than representing signals at a single resolution, INRs can conveniently interpolate signals in a continuous space. This capability is particularly interesting because we can conveniently adjust the required memory making INRs well-suited for high-dimensional 3D reconstructions (Dupont et al., 2021, 2022).
@@
In this paper, we build on INRs and introduce a scalable coordinate-based local reconstruction pipeline for solving imaging inverse problems. In many tasks such as image denoising, super-resolution, and low-dose computed tomography (LDCT), the image intensity at a specific pixel is primarily influenced by the observed image in the pixel’s neighborhood. Our proposed model, termed LoFi (Local Field), recovers the image intensity at each pixel separately, using local information extracted from the input image around that pixel. This information is processed by a neural network (composed of MLP blocks) to determine the image intensity at the target pixel, enabling image reconstruction at any resolution or arbitrary continuous coordinate.
@@
LoFi’s coordinate-based design brings several advantages. Similar to INRs, we can train LoFi on mini-batches of both objects and pixels, resulting in resolution-agnostic memory usage. As a result, LoFi can efficiently process high-resolution images with a tiny memory footprint. Notably, it requires just 3GB of memory to train on 1024×1024 images. Furthermore, LoFi’s local design provides a strong inductive bias for image reconstruction, as demonstrated in our experiments, where LoFi performs effectively on tiny datasets with very few samples, without overfitting or needing regularization and early stopping.
@@
In addition, Deep learning models for image reconstruction like multiscale CNNs (Ronneberger et al., 2015; Zhang et al., 2021) and vision transformers (ViTs) (Liang et al., 2021; Wang et al., 2022; Fabian et al., 2022) often use sophisticated architectures that complicate downstream interpretations. Through our proposed coordinate-conditioned patch geometry, LoFi can learn the position of the relevant features in the input image for each pixel, providing insights for downstream image analysis and interpretation.
@@ II Background and related works
where ${\mathbf{n}}$ is noise. Inverse problems are often ill-posed, implying the existence of multiple plausible reconstructions for a given set of measurements. It is thus important to incorporate an image prior to enable stable and accurate reconstruction. While traditional hand-crafted priors such as total-variation (TV) (Osher et al., 2005) can yield good reconstructions, learning-based priors have brought about substantial improvements (Kothari et al., 2021; Jalal et al., 2021; Chung et al., 2022).
@@ II-A Inverse problems with local operators
Image denoising:
The most straightforward inverse problem is image denoising, where the forward operator ${\mathbf{A}}$ in (1) is the identity matrix; each pixel in the clean image is solely connected to the corresponding pixel in the noisy image, assuming independent noise (between pixels) and natural images. Despite this simplicity, the presence of noise prompts the incorporation of ‘contextual’ information from neighboring noisy pixels. In principle, the required local neighborhood size depends on the noise level and the prior distribution; stronger noises often necessitate larger receptive fields for natural images (Levin and Nadler, 2011).
Low-dose Computed Tomography:
Low-dose computed tomography (LDCT) is a medical imaging technique that reduces radiation exposure to patients during a CT scan by lowering the X-ray beam intensity. Traditional CT imaging uses X-rays to capture cross-sectional images of the body, but these scans can result in significant radiation doses, potentially increasing the risk of long-term side effects like cancer (Brenner and Hall, 2007). LDCT mitigates these risks by significantly lowering the X-ray dose without compromising diagnostic quality, making it a safer imaging process.
@@
${\mathbf{y}}=\mathcal{R}({\mathbf{x}})+{\mathbf{n}},$ (2)
where the forward operator $\mathcal{R}$ is the radon transform applied in angles $\{\alpha_{i}\}_{i=1}^{r}$ and ${\mathbf{n}}$ is the additive noise. In LDCT, the projections ${\mathbf{y}}$ are collected from sufficient angles $r$ but using low-intensity beams resulting in noisy projections. The standard approach for CT image reconstruction is filtered backprojection (FBP),
${\mathbf{x}}_{\text{FBP}}=\mathcal{R}_{\text{FBP}}({\mathbf{y}}).$ (3)
In an ideal scenario where we have an infinite number of noise-free projections, FBP is the exact inverse, namely, ${\mathbf{x}}_{\text{FBP}}={\mathbf{x}}$. In LDCT, we have a sufficient number of noisy projections $N$, the image reconstruction from FBP can be seen as a denoising problem,
${\mathbf{x}}_{\text{FBP}}\approx{\mathbf{x}}+\mathcal{R}_{\text{FBP}}({\mathbf{n}}).$ (4)
The image reconstruction from FBP can be seen as an inverse problem with a local forward operator for LDCT.
Dark matter mapping in cosmology:
As another example, we consider dark matter mapping (Kaiser and Squires, 1993) in cosmology which is important for understanding the nature of dark matter in the Universe through the use of weak gravitational lensing (Mandelbaum, 2018). We will study the convergence that traces the projected mass distribution of dark matter in the Universe. The goal is to estimate the convergence field $\kappa$, the target image, from a related observable, the shear field $\gamma$. The convergence is associated with the magnification of observed galaxies, while the shear corresponds to the image’s stretching due to ellipticity changes. The two fields are related as
@@
where ${\mathbf{n}}$ is the shear field noise, and the forward operator ${\mathbf{A}}$ is a convolutional filter (a Fourier multiplier) with the 2D Fourier transform given by [^1] This is known as the weak lensing planar forward model.,
@@
A simple approach to recover the convergence field $\kappa$ from the shear field $\gamma$ is the Kaiser-Squires (KS) method (Kaiser and Squires, 1993), the current standard in cosmology. The method consists of directly applying ${\mathbf{A}}^{-1}$, the inverse of the forward operator, to the shear field, with a subsequent Gaussian smoothing with an ad hoc setting of the smoothing scale. We consider the naive KS inversion, without the Gaussian smoothing step, as follows,
@@
where $\tilde{{\mathbf{n}}}={\mathbf{A}}^{-1}{\mathbf{A}}{\mathbf{n}}$. The recovery of the convergence field from the naive inverted KS image is thus an inverse problem with a local forward operator.
@@ II-B Deep learning for image reconstruction
Over the last decade, various deep learning approaches have achieved state-of-the-art (SOTA) performance in image reconstruction. Some early approaches segmented the image into patches and employed an MLP to denoise each patch individually (Burger et al., 2012). However, the optimal patch size for high-quality reconstructions depends on the noise level and data distribution. Zhang et al. (Zhang et al., 2017a) introduced DnCNN, an image-to-image CNN designed for image denoising where the network depth is adjusted based on the required receptive field size suggested by traditional methods (Zoran and Weiss, 2011). To expand CNN’s receptive field and use more contextual information, they used dilated convolutions (Zhang et al., 2017b). Concurrently, the U-Net (Ronneberger et al., 2015) has been the backbone model for SOTA image reconstruction architectures (Liu et al., 2018; Park et al., 2019; Jia et al., 2021; Liu et al., 2022), in particular, the DRU-Net (Zhang et al., 2021) achieved remarkable performance by integrating residual blocks (He et al., 2016) into the U-Net architecture. Despite their success, these CNNs lack a systematic procedure for automatic adjustment of the required receptive field size; a tiny receptive field hinders the model performance leading to sub-optimal reconstructions while an excessively large receptive field can result in overfitting the large-scale features and poor generalization.
More recently, vision transformers (ViTs) (Dosovitskiy et al., 2020) have achieved comparable or even better performance than CNNs for various image reconstruction tasks (Chen et al., 2021a; Liang et al., 2021; Wang et al., 2022; Mansour et al., 2022). The core concept of ViTs is to divide the input image into non-overlapping patches, transform each patch into an embedding using a learned linear projection, and then process the resulting tokens through a series of stacked self-attention layers and MLP blocks. The advantage of ViTs over CNNs is attributed to their ability to capture long-range dependencies in the images through the self-attention mechanism (Zamir et al., 2022). A major challenge for ViTs in image reconstruction is the quadratic computational complexity of global self-attention relative to the input image size (Wang et al., 2021). Some approaches have addressed this by designing hierarchical transformers (Heo et al., 2021; Wang et al., 2021). The authors of (Fabian et al., 2022) use convolutional blocks to reduce the spatial dimensions followed by self-attention blocks applied in different scales. Despite the marginally better reconstructions compared to CNNs for inverse problems with non-local forward operators like magnetic resonance imaging (MRI), they show comparable performance with CNNs for local problems like image denoising (Fabian et al., 2022). Moreover, unlike CNNs, ViTs do not benefit from the appropriate inductive bias for image generation, necessitating huge datasets for training (Dosovitskiy et al., 2020; Mansour et al., 2022).
@@
II-C Patch-based image reconstruction
While traditional methods for image reconstruction often process small patches extracted from the input image (Buades et al., 2005; Dabov et al., 2007), deep learning models with patch-based design have recently gained considerable attention. ViTs (Dosovitskiy et al., 2020) and MLP-mixers (Tolstikhin et al., 2021) are a well-known family of patch-based models whose strong performance is often attributed to patch-based design (Trockman and Kolter, 2022). Deep generative models trained on small patches have been used for high dimensional image generation (Ding et al., 2024) and more efficient training on datasets with significantly fewer images (Wang et al., 2024). Patch priors (Gilton et al., 2019; Altekrüger et al., 2022; Hertrich et al., 2022) as efficient alternatives to image priors have been used for image reconstruction which allows training on small datasets with few images well-suited for scientific tasks where we have access to a limited number of ground truth images.

### v2

# LoFi: Neural Local Fields for Scalable Image Reconstruction

@@ Abstract
Neural fields or implicit neural representations (INRs) have attracted significant attention in computer vision and imaging due to their efficient coordinate-based representation of images and 3D volumes. In this work, we introduce a coordinate-based framework for solving imaging inverse problems, termed LoFi (Local Field). Unlike conventional methods for image reconstruction, LoFi processes local information at each coordinate separately by multi-layer perceptrons (MLPs), recovering the object at that specific coordinate. Similar to INRs, LoFi can recover images at any continuous coordinate, enabling image reconstruction at multiple resolutions. With comparable or better performance than standard deep learning models like convolutional neural networks (CNNs) and vision transformers (ViTs), LoFi achieves excellent generalization to out-of-distribution data with memory usage almost independent of image resolution. Remarkably, training on $1024\times 1024$ images requires less than 200MB of memory—much below standard CNNs and ViTs. Additionally, LoFi’s local design allows it to train on extremely small datasets with 10 samples or fewer, without overfitting and without the need for explicit regularization or early stopping.
@@ I Introduction
Imaging inverse problems are ubiquitous in domains like medicine [1], material science [2], remote sensing [3] and cosmology [4]. Deep learning is often the method of choice for solving inverse problems. In particular, convolutional neural networks (CNNs) like U-Net [5] have shown remarkable performance across diverse tasks including computed tomography (CT) [6], magnetic resonance imaging (MRI) [7], radio interferometric imaging [8] and dark matter mapping [9] in cosmology. The success of the U-Net is primarily attributed to its multiscale architecture with a large receptive field that extracts features from the input image at different scales [10].
@@
Despite their strong performance on low-dimensional 2D imaging problems, deep learning architectures can become computationally expensive for large images [11, 12, 13, 14]. This inefficiency arises because current deep neural networks reconstruct the entire image simultaneously, necessitating substantial memory for back-propagation during training, especially for large images. Additionally, interpreting such complex architectures [15, 12, 16] and analyzing their reconstructions can be challenging. Simplified neural architectures can enhance our understanding of reconstruction mechanisms, enabling the design of robust models with strong generalization [17].
@@
Recently, implicit neural representations (INRs) [18, 19, 20] have emerged as a promising coordinate-based pipeline for representing continuous signals, images, and 3D volumes. Unlike most existing deep learning models that treat signals as discrete arrays, INRs map coordinates to signal values using a deep neural network, typically a multi-layer perceptron (MLP), resulting in a continuous signal representation. INRs have several advantages over standard deep learning models. Rather than representing signals at a single resolution, INRs can conveniently interpolate signals in a continuous space. This coordinate-based representation is particularly interesting because we can adjust the required memory making INRs well-suited for high-dimensional 3D reconstructions [21, 22, 23, 24, 25, 26] and scene representation [27].
@@
In this paper, we introduce a scalable coordinate-based local reconstruction pipeline for solving imaging inverse problems. In many tasks such as image denoising and low-dose computed tomography (LDCT), the image intensity at a specific pixel is primarily influenced by the observed image in the pixel’s neighborhood [28]. Our proposed model, termed LoFi (Local Field), recovers the image intensity at each pixel separately, using local information extracted from the input image around that pixel. This information is processed by a neural network (composed of MLP blocks) to determine the image intensity at the target pixel, enabling image reconstruction at any resolution or arbitrary continuous coordinate.
@@
LoFi’s coordinate-based design brings several advantages. Unlike standard CNNs and ViTs which requires substantial memory during training, we can train LoFi, Similar to INRs, on mini-batches of both objects and pixels. This results in resolution-agnostic memory usage, making it highly efficient for high-resolution image processing. As illustrated in Figure 1, LoFi demonstrates significantly lower memory and time requirements compared to CNNs and ViTs. For instance, training LoFi on $1024\times 1024$ images requires less than 200MB memory. Furthermore, its local design provides a strong inductive bias for image reconstruction; unlike CNNs and ViTs, which tend to overfit when trained on tiny datasets with very few samples, LoFi consistently shows robust performance without overfitting, eliminating the need for regularization or early stopping.
@@
In addition, deep learning models for image reconstruction like multiscale CNNs [5, 15] and vision transformers (ViTs) [12, 29, 16, 30] often use sophisticated architectures that complicate downstream interpretations. Through our proposed coordinate-conditioned patch geometry, LoFi can learn the position of the relevant features in the input image for each pixel, providing insights for downstream image analysis and interpretation.
(a) Memory footprint (batch size 1)
(b) Training time (100 iterations)
Figure 1: The memory and time requirements during training for different models; LoFi is significantly faster and more memory-efficient than CNNs and ViTs. Notably, LoFi’s memory usage remains almost independent of image resolution, making it an ideal choice for high-dimensional image reconstruction. All experiments are conducted using a single A100 GPU with 80GB memory. Missing data points indicate that the corresponding model exceeds the GPU memory capacity for the given resolution.
@@ II Background and related works
where ${\mathbf{n}}$ is noise. Inverse problems are often ill-posed, implying the existence of multiple plausible reconstructions for a given set of measurements. It is thus important to incorporate an image prior to enable stable and accurate reconstruction. While traditional hand-crafted priors and regularizers such as total variation (TV) [31] can yield good reconstructions, learning-based priors have brought about substantial improvements [32, 33, 34, 35].
@@ II-A Inverse problems with local operators
Image denoising
The most straightforward inverse problem is image denoising, where the forward operator ${\mathbf{A}}$ in (1) is the identity matrix; each pixel in the clean image is solely connected to the corresponding pixel in the noisy image, assuming independent noise (between pixels) and natural images [28]. Despite this simplicity, the presence of noise prompts the incorporation of ‘contextual’ information from neighboring noisy pixels. In principle, the required local neighborhood size depends on the noise level and the prior distribution; stronger noise usually calls for larger receptive fields [36, 28].
Low-dose computed tomography
Low-dose computed tomography (LDCT) is a medical imaging technique that reduces radiation exposure to patients during a CT scan by lowering the X-ray beam intensity [37]. Traditional CT imaging uses X-rays to capture cross-sectional images of the body, but these scans can result in significant radiation doses, potentially increasing the risk of long-term side effects like cancer [38, 39]. LDCT mitigates these risks by significantly lowering the X-ray dose without compromising diagnostic quality, making it a safer imaging process.
@@
${\mathbf{q}}=\mathcal{R}({\mathbf{f}})+{\mathbf{n}},$ (2)
where the forward operator $\mathcal{R}$ is the radon transform [40] applied in angles $\{\alpha_{i}\}_{i=1}^{r}$ and ${\mathbf{n}}$ is the additive noise. In LDCT, the projections ${\mathbf{y}}$ are collected from sufficient angles $r$ but using low-intensity beams resulting in noisy projections. The standard approach for CT image reconstruction is filtered backprojection (FBP),
${\mathbf{f}}_{\text{FBP}}=\mathcal{R}_{\text{FBP}}({\mathbf{q}}).$ (3)
In an ideal scenario where we have an infinite number of noise-free projections, FBP is the exact inverse, namely, ${\mathbf{f}}_{\text{FBP}}={\mathbf{f}}$. In LDCT, we have a sufficient number of noisy projections $N$, the image reconstruction from FBP can be seen as a denoising problem,
${\mathbf{f}}_{\text{FBP}}\approx{\mathbf{f}}+\mathcal{R}_{\text{FBP}}({\mathbf{n}}).$ (4)
The image reconstruction from FBP can be seen as an inverse problem with a local forward operator for LDCT. For a comprehensive evaluation of current LDCT image reconstruction methods, we refer to [41], where the authors conducted an in-depth analysis of the performance of CNNs, using both simulated and real experimental data for training and inference.
Dark matter mapping in cosmology
As another example, we consider mapping the dark matter distribution of the Universe [4] by solving a weak gravitational lensing [42] inverse problem. We will study the convergence that traces the projected mass distribution of dark matter in the Universe. The goal is to estimate the target convergence field $\kappa$ that traces the projected mass distribution, from a related observable, the shear field $\gamma$. Higher order statistical properties of the convergence field are highly informative for studying the nature of dark energy and dark matter. The two fields are related as,
@@
where ${\mathbf{n}}$ is the shear field noise, and the forward operator ${\mathbf{A}}$ is a convolutional filter (a Fourier multiplier) with the 2D Fourier transform given by, [^1] This is known as the weak lensing planar forward model. Note that ${\mathbf{D}}_{k_{x},k_{y}}$ is undefined for $(k_{x},k_{y})=(0,0)$, meaning that the observable $\gamma$ cannot constrain the average of $\kappa$. This is known as the mass sheet degeneracy. The choice ${\mathbf{D}}_{0,0,}=1$ is made for convenience and is secondary to what follows.
@@
A simple approach to recover the convergence field $\kappa$ from the shear field $\gamma$ is the Kaiser-Squires (KS) method [4], the current standard in cosmology. The method consists of directly applying ${\mathbf{A}}^{-1}$, the inverse of the forward operator, to the shear field, followed by a subsequent Gaussian smoothing with an ad hoc smoothing scale. We consider the naive KS inversion, without the Gaussian smoothing step, as follows,
@@
where $\tilde{{\mathbf{n}}}={\mathbf{A}}^{-1}{\mathbf{A}}{\mathbf{n}}$. The recovery of the convergence field from the naive inverted KS image is thus an inverse problem with a local forward operator. We note that the processed noise $\tilde{{\mathbf{n}}}$ is now spatially correlated and modeling it may still benefit from a larger receptive field; this is addressed by the learnable noise suppression filter described in Section III-D.
@@ II-B Deep learning for image reconstruction
Over the last decade, various deep learning approaches have achieved state-of-the-art (SOTA) performance in image reconstruction. Some early approaches segmented the image into patches and employed an MLP to denoise each patch individually [43]. However, the optimal patch size for high-quality reconstructions depends on the noise level and data distribution [28]. Zhang et al. [11] introduced DnCNN, an image-to-image CNN designed for image denoising where the network depth is adjusted based on the required receptive field size suggested by traditional methods [44]. To expand CNN’s receptive field and use more contextual information, they used dilated convolutions [45]. Concurrently, the U-Net [5] has been the backbone model for SOTA image reconstruction architectures [46, 47, 48, 10], in particular, the DRU-Net [15] achieved remarkable performance by integrating residual blocks [49] into the U-Net architecture.
More recently, vision transformers (ViTs) [50] have achieved comparable or even better performance than CNNs for various image reconstruction tasks [51, 12, 29, 52]. The core concept of ViTs is to divide the input image into non-overlapping patches, transform each patch into an embedding using a learned linear projection, and then process the resulting tokens through a series of stacked self-attention layers and MLP blocks. The advantage of ViTs over CNNs is attributed to their ability to capture long-range dependencies in the images through the self-attention mechanism [12, 30]. As shown in Figure 1, a major challenge for ViTs in image reconstruction is the quadratic computational complexity relative to the input image size [13, 53] which makes ViTs impractical for high dimensional image reconstruction [14]. Recent works attempted to reduce the computational complexity of vision transformers. Restormer [53] applies the self-attention along the channel dimension to reduce memory footprint. Some approaches design hierarchical transformers [54, 13]. The authors of [16] use convolutional blocks to reduce the spatial dimensions followed by self-attention blocks applied in different scales. Despite the marginally better reconstructions compared to CNNs for inverse problems with non-local forward operators like magnetic resonance imaging (MRI), they show comparable performance with CNNs for local problems like image denoising [16].
@@
II-C Neural fields for inverse problems
INRs, particularly those with sinusoidal activations, efficiently model signals and their spatial derivatives. This is particularly useful for solving partial differential equations (PDEs) [18, 55]. In the domain of continuous super-resolution, INRs have been effectively combined with CNNs. For instance, the authors of [56] proposed conditioning the INR’s generation process on local features extracted by a CNN from the low-resolution image. Closer to our work, the authors of [57] used a similar approach, where local features extracted from a low-resolution image are processed with a lightweight CNN to enable continuous super-resolution.
II-D Patch-based image reconstruction
While traditional methods for image reconstruction often process small patches extracted from the input image [58, 59], deep learning models with patch-based design have recently gained considerable attention. ViTs [50] and MLP-mixers [60] are a well-known family of patch-based models whose strong performance is often attributed to patch-based design [61]. Deep generative models trained on small patches have been used for high dimensional image generation [62] and more efficient training on datasets with significantly fewer images [63]. Patch priors [64, 65, 66, 67] as efficient alternatives to image priors have been used for image reconstruction which allows training on small datasets with few images well-suited for scientific applications where we have access to a limited number of ground truth images.